OpenStudy (anonymous):

.A patio in the shape of a semicircle has a diameter of 21 ft. What is the approximate area of the patio, rounded to the nearest square foot? Use 3.14 to approximate pi.

OpenStudy (anonymous):

ok first do 3.14*10.5 to the second power

OpenStudy (anonymous):

then divide that answer by 2 to get your final answer

OpenStudy (anonymous):

how do i do "to the second power"?

OpenStudy (anonymous):

10.5 times itself

OpenStudy (anonymous):

so is it 544 ft2?

OpenStudy (anonymous):

you should have got 173

OpenStudy (anonymous):

can u show ur work?

OpenStudy (anonymous):

10.5*10.5=110.25 110.25*3.14=346.185 346.185/2=173

OpenStudy (anonymous):

thx can u help me with 1 more?

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

.What is the approximate area of the shaded region? A. 104.86 cm² B. 87.47 cm² C. 27.02 cm² D. 10.54 cm²

OpenStudy (anonymous):